OpenStudy (anonymous):

what is the equation of a line that passes through the point (-1 -2) and is perpendicular to -5x=6y+18?

OpenStudy (ivettef365):

First find the slope by y=mx+b

OpenStudy (ivettef365):

the equation is the same as 6y+18=-5x

OpenStudy (ivettef365):

Therefore 6y=-5x-18 y=-(5/6)x-3

OpenStudy (ivettef365):

The slope is -5/6 to be perpendicular it needs to be opposite reciprocal therefore the slope needs to be 6/5